# Linear Complexity Computation of Code Distance and Minimum Size of Trapping Sets for LDPC Codes with Bounded Treewidth

### v1

@@ IV Computer The Size and Number of the Smallest Trapping Set with $b>0$ in Codes with Bounded Treewidth
We provide the pseudocode for the algorithm that computes the size and the number of the smallest trapping sets with a given $b$ based on the algorithm in Theorem 5, which includes the case of $b=0$, in the appendix. In this section, we successfully apply the algorithm to a larger number of spatially coupled LDPC codes, for which a tree decomposition, or even a path decomposition, can be easily found. To verify the correctness of the algorithm, we apply it to short spatially coupled LDPC codes, where all small trapping sets can be enumerated using an exhaustive search algorithm.
Figure 4 presents a randomly generated spatially coupled LDPC code, where the uncoupled subcode corresponds to a $3\times 4$ matrix. The coupling length is 10, and the coupling width (as defined in [35]) is 2. Let $G$ denote the Tanner graph corresponding to the spatially coupled LDPC code. We can easily obtain a path decomposition with a width equal to the product of the number of check nodes in the uncoupled subcode and the coupling width plus one. This is achieved by placing each check node along with all its neighboring variable nodes into the same bag, as illustrated in Figure 5.
By applying the algorithm in Theorem 5 to the graph $G$, we obtain that the size of the smallest trapping set with $b=0$ and $b=1$ is 4 and 3, respectively, in $G$. Additionally, the number of trapping sets with $b=0$ and $b=1$ is 7 and 10 in $G$ respectively. To verify the correctness of these results, we perform an exhaustive search algorithm on $G$ to identify all trapping sets with $b=0$ and $a<4$, as well as all trapping sets with $b=1$ and $a<3$. The results are listed in Table I, demonstrating the accuracy of our proposed algorithm.
TABLE I: All the $(a,b)$-trapping set with $a<4,b=0$ and $a<3,b=1$
number (4,0)(4,0)TS in $G$ (3,1)(3,1)TS in $G$
1 13,14,16,20 1,6,7
2 13,16,18,22 2,3,6
3 14,18,20,22 2,4,6
4 25,27,28,30 2,4,7
5 29,30,31,32 2,6,12
6 29,31,34,35 4,5,9
7 30,32,34,35 4,5,12
8 5,7,12
9 17,26,28
10 21,23,26
Figure 4: A randomly generated spatially coupled LDPC code, where the uncoupled subcode corresponds to a $3\times 4$ matrix. The coupling length is 10, and the coupling width is 2. The variable nodes have a regular degree of 3. The black regions indicate positions in the parity-check matrix where the entries are 1, while other positions contain 0.
Figure 5: A path decomposition of the Tanner graph corresponding to Fig.4 with a width of 10. This indicates that the treewidth and pathwidth of graph $G$ are at most 10.
@@ VI Conclusion
For a given $b$, we present an algorithm to compute the size and the number of the smallest trapping sets with $b\geq 0$. For a Tanner graph $G$ with treewidth $k$, the algorithm runs in $O(b^{2}2^{2k})$ time, while for a Tanner graph $G$ with pathwidth $k$, it runs in $O(b2^{k})$ time. We have validated the correctness of the algorithm both theoretically and through simulation experiments. Furthermore, this algorithm can also be used to find an $(a,b)$-trapping set with the minimum $a$.
Appendix A Algorithm
We provide the pseudocode for the algorithm that computes the size and the number of the smallest trapping sets with a given $b$ based on the algorithm in Theorem 5, which includes the case of $b=0$.
Given the Tanner graph $G$ and a rooted nice tree decomposition $(T,(B_{t})_{t\in T})$ of $G$, for any node $t\in T$ and a given $b$, the algorithm does not compute all $f_{t}(I,Q,d)$ and $g_{t}(I,Q,d)$ over the entire domain $(2^{B_{t}^{c}},2^{B_{t}^{v}},\{0,1,\dots,b\})$. Instead, the values of $f_{t}(I,Q,d)$ and $g_{t}(I,Q,d)$ are only computed when a trapping set with parameters $(I,Q,d)$ is found in $G_{t}$. In the actual implementation, an array or a linked list $O_{t}$ is used to store those tuples $(I,Q,d)$ for which a corresponding trapping set has been identified. The functions $f_{t}$ and $g_{t}$ can be viewed as arrays or linked lists of the same size as $O_{t}$, indexed by the elements stored in $O_{t}$.
As Algorithm 1, accessing the nodes in the tree decomposition following the order of $\mathcal{T}$ ensures that when computing the functions $f_{t}$ and $g_{t}$ for the current node, the corresponding functions $f_{t^{\prime}}$ and $g_{t^{\prime}}$ for its child nodes have already been computed.
Data: Tanner Graph $G$ and it’s a rooted nice tree decomposition $(T,(B_{t})_{t\in T})$ with root $r$, constant $b$
Result: $f_{r}(\emptyset,\emptyset,b),g_{r}(\emptyset,\emptyset,b)$
1 Sort the nodes in $T$ and denote the sorted sequence as $\mathcal{T}$, such that for any node $t$, its children appear before $t$ in the sequences;
2 foreach $t\in\mathcal{T}$ do
3 if $t$ has a child node then
4     Let $t^{\prime}$ be the child of $t$, then $O_{t^{\prime}}$, $f_{t^{\prime}}$, and $g_{t^{\prime}}$ have already been computed;
5   end if
6 if $t$ has two child nodes then
7     Let $t_{1},t_{2}$ be the child of $t$, then $O_{t_{1}},$ $f_{t_{1}}$, $g_{t_{1}}$, $O_{t_{2}}$, $f_{t_{2}}$, and $g_{t_{2}}$ have already been computed;
8   end if
9
10 switch the type of $t$ do
11    case leaf node do
12       $O_{t}=\emptyset,f_{t}=\emptyset,g_{t}=\emptyset$;
13
14    case introduced variable node do
[…]
16         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
17
18    case forget variable node do
[…]
20         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
21
22    case introduced check node do
[…]
24         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
25
26    case forget check node do
[…]
28         delete $O_{t^{\prime}},f_{t^{\prime}}$ and $g_{t^{\prime}}$;
29
30    case join node do
[…]
32         delete $O_{t_{1}},f_{t_{1}},g_{t_{1}},O_{t_{2}},f_{t_{2}}$ and $g_{t_{2}}$;
33
34      end case
35
36   end switch
37
38 end foreach
Algorithm 1 computes the size and the number of the smallest trapping sets with a given $b\geq 0$ in $G$.
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1
2Let $v=B_{t}\setminus B_{t^{\prime}}$;
3 $O_{t}=O_{t^{\prime}}$; $f_{t}=f_{t^{\prime}};g_{t}=g_{t^{\prime}}$;
4 $I_{v}=\Gamma_{o}(G[v\cup B_{t}^{c}])$ ;
5 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}\setminus\{(\emptyset,\emptyset,0)\}$ do
6 $(I,Q,d)=(I_{t^{\prime}}\oplus I_{v}),Q_{t^{\prime}}\cup\{v\},d_{t^{\prime}})$;
7 $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
8 $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})+1$;
9 $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
10
11 end foreach
[…]
13 $f_{t}(I_{v},\{v\},0)=1,g_{t}(I_{v},\{v\},0)=1$ ;
Algorithm 2 IVN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1 Let $v=B_{t^{\prime}}\setminus B_{t}$;
2 $O_{t}=\emptyset$;
3 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}$ do
4
5 if $v\notin Q_{t^{\prime}}$ then
6    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$
7 else
8    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}}\setminus\{v\},d_{t^{\prime}})$
9   end if
10 if $(I,Q,d)\notin O_{t}$ then
11    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
12    $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
13    $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
14
15 else
16    $r=\min\{f_{t}(I,Q,d),f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\}$ ;
17    $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})}g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
18    $f_{t}(I,Q,d)=r\}$;
19
20   end if
21
22 end foreach
Algorithm 3 FVN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1 Let $c=B_{t}\setminus B_{t^{\prime}}$ ;
2 $O_{t}=\emptyset$;
3 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}$ do
4 if $c$ has an odd number of edges connected to $Q_{t^{\prime}}$ in $G_{t}$ then
5    $(I,Q,d)=(I_{t^{\prime}}\cup\{c\},Q_{t^{\prime}},d_{t^{\prime}})$;
6
7 else
8    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
9
10   end if
11 $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
12 $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
13 $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
14
15 end foreach
Algorithm 4 ICN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$
Result: $O_{t},f_{t},g_{t}$
1 Let $c=B_{t^{\prime}}\setminus B_{t}$ ;
2 $O_{t}=\emptyset$;
3 foreach $(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})\in$ $O_{t^{\prime}}$ do
4 if $c\in I_{t^{\prime}}$ then
5    $(I,Q,d)=(I_{t^{\prime}}\setminus\{c\},Q_{t^{\prime}},d_{t^{\prime}}+1)$;
6
7 else
8    $(I,Q,d)=(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
9
10   end if
11 if $d\leq b$ then
12    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
13    $f_{t}(I,Q,d)=f_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
14    $g_{t}(I,Q,d)=g_{t^{\prime}}(I_{t^{\prime}},Q_{t^{\prime}},d_{t^{\prime}})$;
15
16   end if
17
18 end foreach
Algorithm 5 FCN($O_{t^{\prime}},f_{t^{\prime}},g_{t^{\prime}}$)
Data: $G,(T,(B_{t})_{t\in T}),b,O_{t_{1}},f_{t_{1}},g_{t_{1}},O_{t_{2}},f_{t_{2}},g_{t_{2}}$
Result: $O_{t},f_{t},g_{t}$
1 foreach $(I_{t_{1}},Q_{t_{1}},d_{t_{1}})\in O_{t_{1}}$ do
2 foreach $(I_{t_{2}},Q_{t_{2}},d_{t_{2}})\in O_{t_{2}}$ such that $Q_{t_{1}}=Q_{t_{2}}$ do
3    $(I,Q,d)=(I_{t_{1}}\oplus I_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q_{t_{1}}\cup B_{t}^{c}]),Q_{t_{1}},d_{t_{1}}+d_{t_{2}})$;
4    $f^{\prime}(I,Q,d)=f_{t_{1}}(I_{t_{1}},Q_{t_{1}},d_{t_{1}})+f_{t_{2}}(I_{t_{2}},Q_{t_{2}},d_{t_{2}})-|Q|$;
5    $g^{\prime}(I,Q,d)=g_{t_{1}}(I_{t_{1}},Q_{t_{1}},d_{t_{1}})\times g_{t_{2}}(I_{t_{2}},Q_{t_{2}},d_{t_{2}})$;
6    if $(I,Q,d)\notin O_{t}$ then
7       $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
8       $f_{t}(I,Q,d)=f^{\prime}(I,Q,d)$;
9       $g_{t}(I,Q,d)=g^{\prime}(I_{,}Q,d)$;
10
11    else
12
13       $r=\min\{f_{t}(I,Q,d),f^{\prime}(I,Q,d)\}$;
14       $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f^{\prime}(I,Q,d)}g^{\prime}(I,Q,d)$;
15       $f_{t}(I,Q,d)=r$;
16
17      end if
18
19   end foreach
20
21 end foreach
22foreach $(I_{t_{1}},Q_{t_{1}},d_{t_{1}})\in O_{t_{1}}$ such that $Q_{t_{1}}=\emptyset$  do
23 $(I,Q,d)=(I_{t_{1}},\emptyset,d_{t_{1}})$;
24 if $(I,Q,d)\notin O_{t}$ then
25    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
26    $f_{t}(I,Q,d)=f_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})$;
27    $g_{t}(I,Q,d)=g_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})$;
28
29 else
30
31    $r=\min\{f_{t}(I,Q,d),f_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})\}$;
32    $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})}g_{t_{1}}(I_{t_{1}},\emptyset,d_{t_{1}})$;
33    $f_{t}(I,Q,d)=r$;
34
35   end if
36
37 end foreach
38
39foreach $(I_{t_{2}},Q_{t_{2}},d_{t_{2}})\in O_{t_{2}}$ such that $Q_{t_{2}}=\emptyset$ do
40 $(I,Q,d)=(I_{t_{2}},\emptyset,d_{t_{2}})$;
41 if $(I,Q,d)\notin O_{t}$ then
42    $O_{t}=O_{t}\cup\{(I,Q,d)\}$;
43    $f_{t}(I,Q,d)=f_{t_{2}}(I_{t_{1}},\emptyset,d_{t_{2}})$;
44    $g_{t}(I,Q,d)=g_{t_{2}}(I_{t_{2}},\emptyset,d_{t_{2}})$;
45
46 else
47
48    $r=\min\{f_{t}(I,Q,d),f_{t_{2}}(I_{t_{2}},\emptyset,d_{t_{2}})\}$;
49    $g_{t}(I,Q,d)=\mathbb{1}_{r=f_{t}(I,Q,d)}g_{t}(I,Q,d)+\mathbb{1}_{r=f_{t_{2}}(I_{t_{2}},\emptyset,d_{t_{2}})}g_{t_{2}}(I_{t_{1}},\emptyset,d_{t_{2}})$;
50    $f_{t}(I,Q,d)=r$;
51
52   end if
53
54 end foreach
[…]

### v2

@@ V Simulation Results
In this section, we successfully apply the algorithm to spatially coupled LDPC codes, for which a tree decomposition, or even a path decomposition, can be easily found. To verify the correctness of the algorithm, we apply it to short spatially coupled LDPC codes, where all small trapping sets can be enumerated using an exhaustive search algorithm.
Figure 4 presents a randomly generated spatially coupled LDPC code $G$, where the uncoupled subcode corresponds to a $4\times 5$ matrix. The coupling length is 10, and the coupling width (as defined in [35]) is 3. We can easily obtain a path decomposition with a width equal to the product of the number of check nodes in the uncoupled subcode and the coupling width plus one. This is achieved by placing each variable node along with all its neighboring check nodes into the same bag. Furthermore, a nice rooted tree decomposition can be achieved by adding an additional bag to the structure.
By applying the algorithm in Theorem 1 to the graph $G$, we determine that the size of the smallest $(a,0)$ trapping set is 10, and there is one $(10,0)$ trapping set in $G$. Algorithm 1 identifies a $(10,0)$ trapping set $(v_{12},v_{14},v_{20},v_{21},v_{22},v_{24},v_{25},v_{27},v_{30},v_{34})$, which is consistent with the result obtained via exhaustive search, thereby validating the correctness and effectiveness of the proposed algorithm. Furthermore, on a platform with an AMD Ryzen7 4800H CPU, the algorithm in Theorem 1 executed in 0.57 hours, while the exhaustive search algorithm took 50.13 hours, demonstrating an approximately 88-fold improvement in computational efficiency. As the code length and the minimum value of $a$ increase, our algorithm will continue to outperform the exhaustive search, with a time complexity of $O(n)$ compared to the $O(n^{a})$ complexity of the exhaustive search.
Figure 4: A randomly generated spatially coupled LDPC code. The variable nodes have a regular degree of 3. The black regions indicate positions in the parity-check matrix where the entries are 1, while other positions are 0.
@@ VI Conclusion
For any fixed $b\geq 0$, we have shown that for LDPC codes with bounded treewidth, the minimum size of an $(a,b)$ trapping set, the number of minimum trapping sets, and a corresponding trapping set itself can all be computed in time linear in the code length.